OpenStudy (anonymous):

can somebody help explain how to evaluate for this logarithm? 10^(log12-log3)

OpenStudy (anonymous):

i got up to 10^(log4) and i got stuck. i do not know what to do next?

OpenStudy (anonymous):

@fallenangelorchid

OpenStudy (fallenangelorchid):

>.< I'm better with English, history, and writing sorry.

OpenStudy (anonymous):

@phi am i forgetting a property about logarithms? i can't think of anything to do with this

OpenStudy (phi):

the way you "undo" logs is make them the exponent of their base assuming log 3 means log base 10 (we is probably true) then \[ 10^{\log 3} = 3\]

OpenStudy (anonymous):

why is it equal to 3?

OpenStudy (phi):

you should try to remember this rule 10 * 100 = 1000 or , using exponents \[ 10^2 \cdot 10^2= 10^3 \] to remember you add the exponents

OpenStudy (anonymous):

yeah, i understand that. makes sense

OpenStudy (phi):

***why is it 3*** that was an example. \[ 10^{\log (\text{stuff})} = stuff\]

OpenStudy (anonymous):

oh i didn't know that. so it's a rule for when logarithms are found in the exponent of a base?

OpenStudy (phi):

so with \[ 10^{\log12-\log3} \] use the first rule (backwards) to write it as \[ 10^{\log12} \cdot 10^{-\log3} \]

OpenStudy (phi):

in other words we are using \[ 10^{\log12} \cdot 10^{-\log3} = 10^{\log12 + (-\log3)} \]

OpenStudy (phi):

which is the same rule as \[10^1 \cdot 10^2= 10^3 \] (I see I have a typo up above) but with more complicated things than 1 and 2 for the exponent

OpenStudy (anonymous):

ok im following well

OpenStudy (phi):

now what is \[ 10^{\log12} \] ? (see the stuff rule)

OpenStudy (anonymous):

12

OpenStudy (anonymous):

oh cause you add the exponents if they ahve same base right?

OpenStudy (phi):

so we have \[ 10^{\log12-\log3} = 10^{\log12} \cdot 10^{-\log3} \\ = 12 \cdot 10^{-\log3} \] * add the exponents if they ahve same base right?** yes, exactly

OpenStudy (phi):

one more rule \[ a^{-b} = \frac{1}{a^b} \] use that to rewrite \[ 10^{-\log3} \]

OpenStudy (anonymous):

wouldn't it be 10^9 if you add exponents

OpenStudy (phi):

if it was \(10^{12} \cdot 10^{-3} \), but it isn't

OpenStudy (anonymous):

how to add the logs and you got 12?

OpenStudy (phi):

don't lose the bubble... think of log(12) as a number (it is a number, and your calculator will give you a decimal value for it) using the add exponent rule (in reverse) we can say \[ 10^{\log12-\log3} = 10^{\log12} \cdot 10^{-\log3} \\ = 12 \cdot 10^{-\log3} \] now use this rule a^{-b} = \frac{1}{a^b} to rewrite the second term

OpenStudy (phi):

now use this rule\[ a^{-b} = \frac{1}{a^b} \]to rewrite the second term

OpenStudy (anonymous):

oh cause 10^(log12) = 12 and 10^(-log3) which is 1/(10^(log3) so 12*(1/3)?

OpenStudy (phi):

yes

OpenStudy (anonymous):

so 4 is the value? !

OpenStudy (phi):

yes

OpenStudy (anonymous):

wow thank you. my teacher never taught me the 'stuff' rule apparently

OpenStudy (anonymous):

i would fan you again if i could!! xD

OpenStudy (phi):

It's not standard usage but it sticks in my head better.

OpenStudy (anonymous):

one more thing. i have a question on logarithms in general. can i ask you in another?

OpenStudy (phi):

ok

OpenStudy (anonymous):

thanks!